What is the density of air at 75 °F and 1atm pressure? Express in lbm/ft3 and kg/m3.
Answer #1

For imperial value:

T = 75^{\circ}F = (75+460) R = 535\;\;R

and

P = 1\;\;atm=14.7\;\;psia

Now, for air;

R = 0.3704\;\;psia.ft^3/lbm.R

Now, by ideal gas equation;

P = \left ( \frac{m}{V} \right ) RT

\Rightarrow P = \rho RT

\Rightarrow \rho = \frac{P}{RT}

\Rightarrow \rho = \frac{14.7}{0.3704(535)}

\Rightarrow \rho =0.074181\;\;lbm/ft^3

...(Answer)

For SI unit:

T = 75^oF = \left [(75-32)\times \frac{5}{9} \right ]^oC= 23.889^oC = (23.889+273)K

\Rightarrow T =296.889\;\;K

and

P = 1\;\;atm = 101.325\;\;kPa

Now, for air;

R = 0.287\;\;kJ/kg.K

Density will be;

\rho = \frac{P}{RT}

\Rightarrow \rho = \frac{101.325}{0.287(296.889)}

\Rightarrow \rho = 1.189\;\;kg/m^3

...(Answer)